OpenStudy (anonymous):

What number must you add to the polynomial below to complete the square. x^2-6x+1=17 Is it -2,8?

OpenStudy (lncognlto):

9, isn't it? -6/2=-3 -3^2=9.

OpenStudy (anonymous):

It doesn't show in the Choice box.

OpenStudy (anonymous):

http://prntscr.com/1ftj2r

OpenStudy (lncognlto):

Yep, 8, -2 is correct. :)

OpenStudy (lncognlto):

x^2-6x+1=17 x^2-6x =16 x^2-6x + 9 = 16 + 9 (x - 3)^2 = 25 x-3 = + or - 5 x = 8, -2